# HoneySat: A Network-based Satellite Honeypot Framework

### v2

@@ Abstract
Satellites are the backbone of several mission-critical services that enable our modern society to function, for example, GPS. For years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity. However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking interest in satellite security. Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, hurting the generation of security intelligence that can lead to the development of effective countermeasures.
In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat). To provide evidence of HoneySat’s effectiveness, we surveyed experienced SmallSat operators in charge of in-orbit satellites and deployed HoneySat over the Internet to entice adversaries.
Our results show that 90% of satellite operators agreed that HoneySat provides a realistic and engaging simulation of a SmallSat mission. Additionally, HoneySat successfully deceived human adversaries in the wild and collected 22 real-world satellite-specific adversarial interactions. Finally, in a major demonstration of HoneySat’s robustness, we collaborated with an aerospace company to perform a hardware-in-the-loop operation that resulted in HoneySat successfully communicating with an in-orbit, operational SmallSat mission.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
@@ I Introduction
Satellites are complex devices designed to withstand outer space conditions. They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [1], Earth observation, e.g., the Sentinel constellation [2], and broadband internet service, e.g., the Starlink constellation [3]. In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [4]. As a result, the software and hardware components that make up a specific spacecraft vary greatly. Likewise, satellite missions vary a lot depending on the owner’s budget. For example, university missions are smaller as they have a limited budget [5]. A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend. Such an attack could lead to the cessation of air traffic and widespread communication blackouts [6]. It could also cause food shortages and the freezing of financial transactions [7]. Furthermore, such an attack could exacerbate the Kessler Syndrome [8], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
@@
In this increasingly vulnerable environment, the probability of a successful satellite cyberattack continues to rise. This is driven by three key trends [6]: first, satellite deployments have increased at an unprecedented pace. For instance, while an average of 82 launches took place between 2008 and 2017, as many as 197 launches occurred in 2023 alone, each typically carrying multiple satellites [9]. This surge is partly fueled by the rise of cheaper commercial off-the-shelf (COTS) components and the availability of more Space Launch Vehicles (SLVs), which makes access to orbit affordable for smaller institutions, such as universities [10]. Second, ground station technology has become significantly more affordable (and sometimes open source), greatly lowering the communication barrier with satellites [11]. Thus, a broader range of malicious actors can now communicate with satellites. Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [12].
@@
Although satellite security research has gained increased attention [12, 13], the relationship between outer space and cyberspace remains poorly understood [14]. At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying the tactics, techniques, and procedures (TTPs) used against these critical systems are limited. MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [15], highlighting a significant data gap. Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [7]. As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
A honeypot’s ability to collect real-world cyberattack data makes it an ideal solution to this problem [16]. A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [17]; all the while, the honeypot logs all the interactions the attackers make. This log data can later be analyzed to discover new and existing TTPs.
Since the release of the first honeypot, the Deception Toolkit, in 1997 [18], a wide range of honeypots with ever-increasing capabilities have been introduced. These honeypots are used by universities, companies, and nation-states worldwide [19, 20, 21, 22, 23]. Honeypots are also used to deter malicious actors from attacking different types of systems, from industrial control systems [23] to social media platforms [24]. However, as of the time of writing, there is no space-sector specific honeypot in the literature.
In this paper, we design, implement, and evaluate the first satellite honeypot, HoneySat, to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [25, 26, 27, 28]. HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and satellite). Specifically, HoneySat simulates Small Satellites or SmallSats, which are spacecraft with a mass of less than 180 kilograms [29]. CubeSats, for example, are SmallSats. Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data. This capability enables the creation of complete testing environments for satellite software that integrate a variety of simulated subsystems and sensors. For instance, Satellite Simulator can simulate orbital mechanics (e.g., position tracking, attitude adjustment) and electrical power systems (e.g., power generation, consumption, and distribution) alongside other subsystems.
@@
Moreover, in a significant showcase of HoneySat’s practical relevance, a leading multinational space agency has integrated HoneySat into operational test environments to evaluate its effectiveness in securing mission-critical space systems.
@@ II-A Types of Honeypots
Low-Interaction Honeypots. These honeypots offer minimal interaction, simulating real systems through scripts. They are easy to setup and maintain due to low resource consumption, and have a reduced risk of adversarial takeover. However, they provide limited interaction opportunities to adversaries, which limits the data they provide. Low-interaction honeypot examples include Conpot [30] and Honeyd [31].
High-Interaction Honeypots. These honeypots offer extensive interaction opportunities via emulation or advanced simulations [31]. Their main advantage is providing adversaries with almost limitless interactions, providing extensive interaction data. However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [32]. High-interaction honeypot examples include Cowrie [33] and HoneyPLC [23].
@@ II-B Honeypot’s State of the Art
The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [34, 35]. From implementations that simulate a host’s TCP/IP stack, such as Honeyd [31], to modern approaches that integrate social media applications, such as HoneyTweet [24]. In the absence of a satellite honeypot, we examine the approaches most related to satellites: Industrial Control Systems (ICS) and Unmanned Aerial Vehicles (UAV), summarized in Table I.
Satellite systems like ICS must be aware of some physical process, e.g., the sun’s position, via sensors to acquire data about the physical world. Several ICS honeypots have simulated these physical processes. For example, ICSnet [36] and HoneyICS [37] simulate components such as programmable logic controllers (PLCs) and actuators such as valves.
HoneyDrone [38] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios. Although UAV honeypots share some similarities with satellites [39], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec. II-C.
@@
Physics
Sims
Extensi-
bility
Conpot [30]
@@ II-C Anatomy of a Satellite Mission
Ground Segment 1. The Ground Segment (GS) covers the terrestrial infrastructure required for a successful satellite operation. It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [12]. The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data.
Space Segment 2. The space segment comprises a satellite or a constellation of satellites. A satellite is launched into orbit and then establishes communications with the ground segment. During regular operations, satellites may communicate through one or multiple ground stations [12].
Telecommands (TC) 3 and Telemetry (TM) 4. The basic data flow between the space and ground segments are TC and TM [41]. TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [41]. TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec. II-E). The design and implementation of TCs varies depending on the satellite mission. From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [12].
Orbital Pass 5. A satellite and its ground station can communicate only during an orbital pass. An orbital pass, or more commonly a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication. A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects between the satellite and the ground station, e.g., mountains [42]. In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly. The occurrence of passes can be calculated using the spacecraft’s orbital elements in the two-line element (TLE) data [43].
Satellite Mission Operations 6. Satellite mission operations vary widely depending on the owning organization’s budget, and technology [5]. Nevertheless, they share some commonalities, which we now describe.
A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [44]. Operations are carried out in a Mission Operations Center (MOC), where operators sit at their workstations to manage TCs sent to the satellite(s). Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [11].
@@
Figure 1: The components commonly found in a satellite mission in the context of the space and ground segments.
Ground (Segment) Software (GSS). GSS allows operators to carry out the mission’s routine operations. GSS are very diverse. Some satellite missions develop their own GSS while others use open-source [45] or proprietary GSS [46]. There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
Mission control software 8 manages TCs and scripts to be sent to the satellite and display TM. For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [47]. Some missions develop their own MCS. For example, the SUCHAI mission developed their own MCS application [45].
Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide information about each satellite’s pass. For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [48].
@@ II-D Satellite Architecture
Satellite architectures are varied and complex [49]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s activities and the payload. The platform provides the possibility to run a payload that fulfills the purpose of these missions. Payloads are different depending on the satellite’s mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
Platform: The platform is composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations. These include the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e. attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
These subsystems are controlled via TCs sent to the satellite. TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [50]. This managerial duty is handled by the Flight Software (FS) running on an embedded system. Some examples include NASA’s Core Flight System (cFS) [51] and the University of Chile’s SUCHAI FS [52]. Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all spacecraft subsystems.
Payload. The payload is the equipment that a satellite employs to fulfill its mission. Due to satellites’ varied missions, payloads are heavily customized [53]. For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [54].
@@ II-E Small Satellite Protocol Ecosystems
Cubesat Space Protocol. The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [55]. CSP is implemented as an open-source C library called libCSP [55] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [56].
@@
In case the satellite is not currently performing a pass, the packet is dropped. There are also default services running on standardized ports that are enabled by default when building libCSP [55], for example for network diagnostic (ping) or basic operations (rebooting).
The straightforwardness of CSP makes it particularly relevant for the honeypot use case as the internal structures of any CSP-based mission will look similar, so distinguishing two missions may be difficult.
CCSDS Space Communication Protocols. The Consultative Committee for Space Data Systems (CCSDS) Space Communication protocols are a set of standardized protocols used for different purposes in space communications [57]. A relevant CCSDS protocol for SmallSat TM and TC is the Space Packet protocol. This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [58] to define how TCs and TM are encoded and transported. PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [57].
Proprietary Protocols. In addition to CSP and CCSDS, vendor-specific protocols may be used on top of more standard protocols. Some basic small satellites may have entirely ad-hoc protocols used on a single mission, for example the ESTCube-1 uses a custom protocol [59]. These proprietary protocols are not ideal candidates for constructing honeypots as the ecosystem of one mission will look out of place for other missions, preventing generalization.
The choice of ecosystem also influences the choice of GSS as an MCS usually supports a single protocol stack [45]. In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services. These protocols include Telnet, SSH, VNC, FTP, and HTTP. Telnet, SSH, and VNC are used to remotely connect to the MOC workstations [11]. HTTP is used for web interfaces that operators use to visualize mission data. FTP is used to download TM data from the mission data repository. If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
@@ II-F Space Systems’ Tactics, Techniques and Procedures (TTPs)
Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [60, 61]. Several frameworks have been introduced to standardize space systems’ TTPs. There are two MITRE-style frameworks, the SPARTA matrix [62] and the SPACE-SHIELD matrix [63]. These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise. For the remainder of this paper we use the SPACE-SHIELD matrix to analyze HoneySat.

### v3

@@ Abstract
Satellites are the backbone of mission-critical services that enable our modern society to function, for example, GPS. For years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity. However, technological advancements have made these assumptions obsolete, paving the way for potential attacks. Unfortunately, there is no way to collect data on satellite adversarial techniques, hindering the generation of intelligence that leads to the development of countermeasures.
In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat). To provide evidence of HoneySat’s effectiveness, we surveyed SmallSat operators and deployed HoneySat over the Internet.
Our results show that 90% of satellite operators agreed that HoneySat provides a realistic simulation. Additionally, HoneySat successfully deceived adversaries in the wild and collected 22 real-world adversarial interactions. Finally, we performed a hardware-in-the-loop operation where HoneySat successfully communicated with an in-orbit, operational SmallSat mission.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
@@ I Introduction
Satellites are complex devices designed to withstand outer space conditions. They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [space2023global]. In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [uwe1_wuerzburg_2005]. As a result, the software and hardware components that make up a specific spacecraft vary greatly. A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend. Such an attack could lead to the cessation of air traffic and widespread communication blackouts [holmes2024growing]. It could also cause food shortages and the freezing of financial transactions [vancamp2022world]. Furthermore, such an attack could exacerbate the Kessler Syndrome [drmola2018kessler], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
@@
In this increasingly vulnerable environment, the probability of a successful satellite cyberattack continues to rise. This is driven by three key trends [holmes2024growing]: first, satellite deployments have increased at an unprecedented pace. For instance, while an average of 82 launches took place between 2008 and 2017, as many as 197 launches occurred in 2023 alone, each typically carrying multiple satellites [semanik2023private]. Second, ground station technology has become significantly more affordable (and sometimes open source), greatly lowering the communication barrier with satellites [california2024earth]. Thus, a broader range of malicious actors can now communicate with satellites. Third, satellite engineers and operators continue to rely on protocol obscurity practices, such as hiding specialized knowledge about the transmission protocol implemented on satellites [willbold2023space].
@@
Although satellite security research has gained increased attention [willbold2023space, scharnowski2023case], the relationship between outer space and cyberspace remains poorly understood [pavur2022building]. Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [vancamp2022world]. As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
@@
In the spirit of open and reproducible science, HoneySat’s source code and experimental results are available online [^1] https://doi.org/10.5281/zenodo.17871431, [^2] https://github.com/HoneySat.
@@ II-A Types of Honeypots
Low-Interaction Honeypots. These honeypots offer minimal interaction, simulating real systems through scripts. They are easy to setup and maintain, and have a reduced risk of adversarial takeover. However, they provide limited interaction opportunities, which limits the data they provide. Some examples include Conpot [vestegaard2014conpot] and Honeyd [provos2007virtual].
High-Interaction Honeypots. These honeypots offer extensive interaction opportunities via emulation or simulation [provos2007virtual]. Their main advantage is providing adversaries with almost limitless interactions. However, they pose a high takeover risk [marra2024feasibility]. Examples include Cowrie [oosterhof2024cowrie] and HoneyPLC [lopez2022honeyplc].
@@ II-B Honeypot’s State of the Art
The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [ilg2023survey, nazario2024awesome]. From implementations that simulate a host’s TCP/IP stack, such as Honeyd [provos2007virtual], to modern approaches that integrate social media applications, such as HoneyTweet [acharya2024conning]. In the absence of a satellite honeypot, we examine the approaches most related to satellites: Industrial Control Systems (ICS) and Unmanned Aerial Vehicles (UAV), summarized in Table I.
Satellite systems like ICS must be aware of some physical process, e.g., the sun’s position, via sensors to acquire data about the physical world. ICS honeypots, such as HoneyICS [lucchese2023honeyics], simulate these physical processes. In addition to ICS, UAV honeypots such as HoneyDrone [daubert2018honeydrone] achieve this via simulations, and recreate attack scenarios.
@@
Physics Sims
Extensibility
Conpot [vestegaard2014conpot]
@@ II-C Anatomy of a Satellite Mission
Ground Segment 1. The Ground Segment (GS) covers the terrestrial infrastructure required for a successful satellite operation. It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [willbold2023space]. The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data.
Space Segment 2. The space segment comprises a satellite or a constellation of satellites. A satellite is launched into orbit and then establishes communications with the ground segment. During regular operations, satellites may communicate through one or multiple ground stations [willbold2023space].
Telecommands (TC) 3 and Telemetry (TM) 4. The basic data flow between the space and ground segments are TC and TM [esa2013telemetry]. TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [esa2013telemetry]. TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec. II-E). The design and implementation of TCs varies depending on the satellite mission. From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [willbold2023space].
Orbital Pass 5. A satellite and its ground station can communicate only during an orbital pass. An orbital pass, or simply a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication. A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects, e.g., mountains [wood2006introduction]. Passes can be predicted using two-line element (TLE) data [vallado2012two].
Satellite Mission Operations 6. Satellite mission operations vary widely depending on the owning organization’s budget, and technology [maya1_inquirer_2018]. Nevertheless, they share some commonalities, which we now describe.
A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [nasa2024flight]. Operations are carried out in a Mission Operations Center (MOC), where operators sit at their workstations to manage TCs sent to the satellite(s). Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [california2024earth].
@@
Figure 1: The components commonly found in a satellite mission.
Ground (Segment) Software (GSS). GSS allows operators to carry out the mission’s routine operations. GSS are very diverse. Some satellite missions develop their own GSS while others use open-source [suchai-flight-software] or proprietary GSS [gomspace2021nanocom]. There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
Mission control software 8 manages TCs and scripts to be sent to the satellite and display TM. For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [merri2007cutting]. Some missions develop their own MCS. For example, the SUCHAI mission developed their own MCS application [suchai-flight-software].
Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide information about each satellite’s pass. For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [csete2023gpredict].
@@ II-D Satellite Architecture
Satellite architectures are varied and complex [ivancic2003architecture]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s activities and the payload. The platform underpins the payload that fulfills the mission’s purpose. Payloads differ depending on the satellite’s mission and can range from measurement instruments to communications systems.
Platform: The platform is composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations. These include the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e. attitude); the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM.
These subsystems are controlled via TCs sent to the satellite. TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem [yost2023nasa]. This managerial duty is handled by the Flight Software (FS) running on an embedded system, for example, NASA’s Core Flight System (cFS) [mccomas2016core]. Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all subsystems.
Payload. The payload is the equipment that a satellite employs to fulfill its mission. Due to satellites’ varied missions, payloads are heavily customized [esa2018payload]. For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [nasa2019remote].
@@ II-E Small Satellite Protocol Ecosystems
Cubesat Space Protocol. The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [LibcspLibcsp2024]. CSP is implemented as an open-source C library called libCSP [LibcspLibcsp2024] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [GetStarted].
@@
CCSDS Space Communication Protocols. The Consultative Committee for Space Data Systems (CCSDS) Space Communication is a set of standardized protocols used for different purposes in space communications [pusecosysytem]. A relevant CCSDS protocol for SmallSat TM and TC is the Space Packet protocol. This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [kaufeler1994esa] to define how TCs and TM are encoded and transported. PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [pusecosysytem].
@@ II-F Space Systems’ Tactics, Techniques and Procedures (TTPs)
Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [raza2024what, nist2024tactics]. The SPACE-SHIELD matrix [esa2024space-shield] is a framework used to standardize space systems’ TTPs, for example, ground segment compromise.